OpenStudy (anonymous):

Hng--- Can someone /actually/ help me? (7,0) (-4,0) (-13,0) Using those points as zeros construct the polynomial function f(x) that will be the path of your rollercoaster. Show all of your work.

OpenStudy (kc_kennylau):

FACTOR THEOREM: If (x-a) is a factor of f(x), then f(a)=0 CONVERSE OF FACTOR THEOREM: If f(a)=0, then (x-a) is a factor of f(x)

OpenStudy (kc_kennylau):

You'll want to use the converse of factor theorem.

OpenStudy (mathmale):

So: think of what "factor" of a polynomial means. Take those 3 factors and multiply them out. Try this here and now, please.

OpenStudy (mathmale):

Example: (x-1)(x+2) = x^2 + x -2

OpenStudy (anonymous):

f(a)= 7-a f(a)= -4-a f(a)= -13-a

OpenStudy (mathmale):

(x-1)(x+2) = x^2 + x -2 is an example of what I'm asking you to do: multiply out the factors, using the FOIL method of multiplying out binomials.

OpenStudy (anonymous):

(7-1)(7+2)= 49+14-7-2

OpenStudy (mathmale):

(7,0) (-4,0) are 2 of the 3 horizontal intercepts of the graph of the polynomial you're looking for. Using the guidance that Kenny has given you, the corresponding factors of this poly are (x-7)(x+4). Please multiply that out now.

OpenStudy (anonymous):

(x-7)(x+4)= x^2+4x-7x-28

OpenStudy (mathmale):

Very nice. Would you please simplify that by combining any like terms.

OpenStudy (anonymous):

x^2-3x-28

OpenStudy (mathmale):

Great. This is the product of the first 2 factors of your polynomial, the factors based upon (7,0) and (-4,0). The third horizontal intercept is

OpenStudy (mathmale):

(x+13). Multiply out: (x+13)(x^2-3x-28).

OpenStudy (anonymous):

x^3-3x^2-28x+13x^2-39x-364 x^310x^2-67-364

OpenStudy (anonymous):

x^3+10x^2-67-364 @mathmale

OpenStudy (mathmale):

Nice, but wouldn't that be 67x instead of just 67?

OpenStudy (mathmale):

Still here? Do you happen to know the technique of synthetic division?

OpenStudy (anonymous):

Yeah, sorry. Yeah, it's be 67x. Uhm, I do know the technique of synthetic division. But it's kind of scratchy, why?

OpenStudy (mathmale):

Once you've multiplied out your 3 factors to obtain a polynomial, you could use synth div to determine whether any of the original roots (7, -4 or -13) are actually roots of OUR new polynomial.

OpenStudy (mathmale):

I'll leave it to you whether or not you want to try that. I've made several attempts to check the known root -4, but so far haven't gotten confirmation that our polynomial is correct.

OpenStudy (anonymous):

I have no idea, man...

OpenStudy (mathmale):

Another way of checking your answer might be wolframalpha.com. Please go to this link: http://www.wolframalpha.com/input/?i=factor+x%5E3%2B10x%5E2-67x-364 What would you conclude about your polynomial, based upon the information there?

OpenStudy (anonymous):

That it's correct.

OpenStudy (mathmale):

Cool. wolframalpha.com is a powerful tool for checking your results, making quick and accurate graphs, etc., but not a substitute for knowing how to carry out math operations yourself. Congrats. Why not move on to the next problem?

OpenStudy (anonymous):

Sure. c: